# Event Topology Classifiers at the Large Hadron Collider

### v1

@@ Acknowledgments
S.P. and B.S. acknowledge the doctoral fellowships from the UGC and CSIR, Government of India. S.T. acknowledges the CERN Research Fellowship and the funding received from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101149298. The authors acknowledge the DAE-DST, Government of India, funding under the Mega-Science Project—“Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123). The authors acknowledge Neelkamal Mallick for the discussion on the heavy-ion related studies. S.T. acknowledges discussion with Alexander Philipp Kalweit and Alice Ohlson during the preparation of this manuscript.

### v2

@@ Acknowledgments
S.P. and B.S. acknowledge the doctoral fellowships from the UGC and CSIR, Government of India. S.T. acknowledges the CERN Research Fellowship and the funding received from the European Union’s Horizon Europe research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 101149298. The authors acknowledge the DAE-DST, Government of India, funding under the Mega-Science Project—“Indian participation in the ALICE experiment at CERN” bearing Project No. SR/MF/PS-02/2021- IITI (E-37123). The authors acknowledge Neelkamal Mallick for the discussion on the heavy-ion-related studies. S.T. acknowledges discussion with Alexander Philipp Kalweit and Alice Ohlson during the preparation of this manuscript.